OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4.

OpenStudy (anonymous):

slope= 1, intercept= -4/3

OpenStudy (anonymous):

This can be found by the standard equation of a straight line: y=mx+c where, m=slope, c= intercept.

OpenStudy (anonymous):

slope = 3; y intercept at (0, -1)?

OpenStudy (anonymous):

Calculate it. it's <-1

OpenStudy (anonymous):

so it is that? or possitive 1

OpenStudy (anonymous):

No. not positive.

OpenStudy (anonymous):

Well my choices are slope = -1; y intercept at (0, -negative four thirds) slope = 1; y intercept at (0, -negative four thirds) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1)

OpenStudy (rajee_sam):

3x - 3y = 4 3x - 4 = 3y x - 4/3 = y y = x - 4/3 this is of the form y = mx + b m = slope and b = y-intercept m = 1 and y - int = (0, -4/3)

OpenStudy (anonymous):

2nd option. Definitely. Told ya.

OpenStudy (anonymous):

My 'best answer' vote, lass?